hi....so i really need help i just cant understan direct joint and inverse variations....so ok here is the problem.
Q: if y varies directly as x and y=9 when x=6, find y when x=8

Answer by stanbon(48558)

if y varies directly as x and y=9 when x=6, find y when x=8
---------
If y varies directly as x, y = kx
You 1st need to solve for "k":
9 = k*6
k= 3/2
------------
So the proper equation for your problem is
y = (3/2)x
-------------
Then, if x = 8 you can solve for "y":
y = (3/2)*8 = 12
======================
Comment:
If you have indirect variation y = k/x
Joint is the same as direct
inverse is the same as indirect
-----------------
Cheers,
Stan H.